# WiFi-Diffusion: Achieving Fine-Grained WiFi Radio Map Estimation with Ultra-Low Sampling Rate by Diffusion Models

### v1

@@ IV WiFi-Diffusion: Design Details
WiFi-Diffusion comprises three blocks: a boost block, a generation block, and an election block. In this section, we describe design details of each block of WiFi-Diffusion.
@@ IV-A Boost block
The objective of this block is to improve the qualities of radio maps produced by DDIM in the subsequent generation block. In this block, we design AttUnet, a U-Net model with self-attention, to pinpoint “critical points” within the region $\mathcal{Y}$. AttUnet predicts RSS values at these critical points based on prior information including layout of buildings, law of radio propagation, and RSS values of known radio samples. These predictions are then provided to DDIM. While DDIM, as a diffusion model, is adept at focusing on important features, it lacks a deep understanding of the underlying physical laws. By incorporating AttUnet’s predictions that are aware of physical laws, DDIM’s performance can be improved. Furthermore, AttUnet enhances DDIM by providing it with RSS information of additional radio samples (i.e., with RSS values predicted for the identified critical points). By using both the identified critical points and the limited collected samples, DDIM can generate radio maps of higher quality compared to those maps generated solely from the limited collected samples.
It should be clear that DDIM can more easily predict the propagation of high-power radios in an environment without buildings. Conversely, predicting the propagation of low-power radios in a complex environment with buildings poses a greater challenge for DDIM. We consider the complex environment with buildings as the area containing points that are critical to DDIM.
Figure 4: Architecture of AttUnet used in WiFi-Diffusion.
@@
Specifically, AttUnet first generates a rough estimation of the radio map. The architecture of AttUnet is illustrated in Fig. 4, with its parameters detailed in Table II. Then, we consider that critical points are located in areas where there is a localized, abrupt change in the RSS or a sudden variation in the rate of change in the RSS in the map. These phenomena are typically attributed to the intricate interactions between electromagnetic waves and buildings, as well as other waves. To identify these critical points, we shift a local window in various directions to analyze the area within the window. If this shifting results in substantial changes in the intensity of RSS, we consider the corresponding point to be critical.
@@
where $\mathbf{x}_{a}(x,y)$ is the RSS of the point located at $(x,y)$ in the map, and $w(x,y)$ is a Gaussian window function which could effectively remove the noise in the map and is suitable for smoothing the map with continuity due to its characteristics of high center weights and low edge weights. Local quadratic approximation of $E(u,v)$ in the neighborhood of $(0,0)$ is given by the second-order Taylor expansion:
@@
Given a point $(x,y)$, we first calculate $\mathbf{M}$ and then compute $\lambda_{1}$ and $\lambda_{2}$. If $\lambda_{1}$ and $\lambda_{2}$ are large, the RSS change of the map $\mathbf{x}_{a}$ at this point is abrupt and we consider this point as one critical point.
Overall, in the boost block, we first use AttUnet to generate a preliminary radio map from the known radio samples and building layout, and then identify critical points in the map from the perspective of laws of radio propagation. At the end of the boost block, we convey the predicted radio samples (identified critical points as well as corresponding predicted RSS values) and the known samples collected by sensors to DDIM. To further enhance the performance of DDIM, we also randomly select a specific number of points in the map generated by AttUnet and convey their predicted RSS values to DDIM. Similar to $\mathbf{S}_{k}$, we define $\mathbf{ES}_{k}$ as an $H\times W$ matrix, where the element at index $(h_{s},w_{s})$ represents the RSS value at that grid point if it is a critical point, a randomly selected point, or a point where a radio sample has been collected; otherwise, it is 0. Fig. 5a gives an example of $\mathbf{S}_{k}$ and Fig. 5c gives an example of $\mathbf{ES}_{k}$. Fig. 5b gives an example of one radio map, where there are original points corresponding to known radio samples, critical points identified in the boost block, and randomly selected points in the map. $\mathbf{ES}_{k}$ is the output of the boost block and is the input for the subsequent generation block.
Figure 5: Different kinds of points identified by the boost block in WiFi-Diffusion.
@@ IV-B Generation block
The objective of this block is to generate a diverse set of radio map candidates. We employ DDIM to achieve this goal.
We consider the task of radio map estimation as a generative problem, wherein a radio map is regarded as a sample drawn from an $(H\times W)$-dimensional distribution. By learning this distribution from available radio map datasets, we can effectively tackle the challenge posed by ultra-low sampling rate and produce high-quality radio maps even when the number of known radio samples in the map to be estimated is extremely small.
We assume that $\mathbf{x}_{0}$ which is the radio map to be estimated follows a certain conditional probability distribution $\mathcal{X}_{c}$ given $\mathbf{B}$ and $\mathbf{ES}_{k}$. We use Denoising Diffusion Probabilistic Model (DDPM) [20] to learn $\mathcal{X}_{c}$. The key idea of DDPM involves transforming $\mathcal{X}_{c}$ into a standard normal distribution $\mathcal{N}(\mathbf{0},\mathbf{I})$, and then trains a neural network to model the reverse process, thereby establishing a mapping from $\mathcal{N}(\mathbf{0},\mathbf{I})$ back to $\mathcal{X}_{c}$. With DDPM, we can get sample values of $\mathcal{X}_{c}$ by sampling in $\mathcal{N}(\mathbf{0},\mathbf{I})$.
Specifically, DDPM consists of two processes: forward process (diffusion process) and reverse process (inverse diffusion process).
@@
where $\beta_{t}$ is the hyperparameter chosen before model training. Using the chain rule and Markov property of probability, we denote the joint distribution of $\mathbf{x}_{1},\mathbf{x}_{2}\dots\mathbf{x}_{t}$ conditional on $\mathbf{x}_{0}$ as:
@@
where $\epsilon\sim\mathcal{N}(\mathbf{0},\mathbf{I})$. Let $\alpha_{t}:=1-\beta_{t},\bar{\alpha}_{t}:=\prod_{s=0}^{t}\alpha_{s}$, and we can marginalize the above joint distribution to give:
@@
Define the condition $\{\mathbf{B},\mathbf{ES_{K}}\}$ as $c$, we have:
@@
During inference, we aim to generate a diverse set of results (radio maps) from different noises sampled in $\mathcal{N}(\mathbf{0},\mathbf{I})$. A significant limitation of DDPM is the substantial number of iterations required to infer and generate high-quality results. To address this issue, we use DDIM [25] for inference. DDIM markedly reduces the inference time while preserving the high quality of the generated results. Furthermore, DDIM can be seamlessly applied after the model training block, allowing for a straightforward conversion of a DDPM model to a DDIM model without the necessity of retraining.
@@
where $\sigma_{t}^{2}=\eta\cdot\sqrt{(1-\alpha_{t-1})/(1-\alpha_{t})}\sqrt{(1-\alpha_{t}/\alpha_{t-1})}$. If $\eta=1$, the DDIM method is the same as DDPM. If $\eta=0$, the DDIM method is to solve the ordinary differential equations.
The training and inference of the generation block are illustrated in Fig. 6. At the end of the generation block, we get a diverse set of radio map candidates generated by DDIM from $\mathbf{B}$ and $\mathbf{ES}_{k}$.
@@ IV-C Election block
The objective of this block is to identify the best map in the diverse set of radio map candidates generated by the generation block. We leverage the radio propagation model as a guide to identify the map that most closely adheres to the propagation model in the set and consider it as the best map.
@@
As discussed in the previous block and illustrated by Fig. 7, the diffusion model DDIM can build a mapping from the noisy distribution $\mathcal{N}(\mathbf{0},\mathbf{I})$ to the conditional distribution $\mathcal{X}_{c}$ under condition $c$
@@
where, $\mathbf{x}_{zi}\in\mathcal{X}_{c}$, $\mathbf{z}_{i}\in\mathcal{N}(\mathbf{0},\mathbf{I})$, $i\in\{1,2,\cdots,m\}$, and $m$ is the number of map candidates generated by DDIM.
The problem of radio map estimation requires the result $\mathbf{x}_{zi}$ to be as close as possible to the true radio map $\mathbf{x}_{0}$. Therefore, we generate $\{\mathbf{x}_{z1},\mathbf{x}_{z2},\cdots,\mathbf{x}_{zm}\}$ from $\{\mathbf{z}_{1},\mathbf{z}_{2},\cdots,\mathbf{z}_{m}\}$, and select $\mathbf{x}_{z}$ that is closest to $\mathbf{x}_{0}$. Considering that the true map $\mathbf{x}_{0}$ is unknown, intuitively, we plan to use the Mean-Square Error (MSE) loss between $\mathbf{S}_{k}$ consisting of the RSS values of known radio samples and the RSS values predicted by DDIM to approximate the distance between $\mathbf{x}_{zi}$ and $\mathbf{x}_{0}$, i.e.
@@
where $\mathbf{mask}_{k}$ denotes a mask matrix consisting entirely of zeros, except for the element located at index $(h_{s},w_{s})$ which corresponds to a non-zero element in $\mathbf{S}_{k}$. From the above equation we have
@@
However, we note that $\text{spMSE}(\mathbf{x}_{zi},\mathbf{S}_{k})$ does not approximate $\text{MSE}(\mathbf{x}_{zi},\mathbf{x}_{0})$ well in practice, mainly due to the ultra-low sampling rate (which means that almost all elements in $\mathbf{S}_{k}$ are zero). Recall that in the boost block, AttUnet generates a radio map $\mathbf{x}_{a}$. We use $\mathbf{x}_{a}$ to help select the best map. We introduce a correction factor $res$ as follows
@@
where the value of $res$ can be derived from different methods of analyses. ResNet [16] has been widely used over the past few years in feature extraction applications. In WiFi-Diffusion, we use it to predict $res$. We do this by providing $\mathbf{x}_{a}$, $\mathbf{x}_{zi}$, Laplacian of Gaussain (LoG) operator processing information (edge detection), and high power maps of $x_{zi}$ (an example for these information is shown in Fig. 8) to ResNet to learn the correction factor $res$.
@@
In addition to utilizing the MSE as a criterion for selecting the best map during the election block, we also incorporate radio propagation models as a guiding tool. It is important to note that obtaining an exact radio propagation model for the radio map being estimated is almost impossible, particularly in complex environments. Therefore, instead of concentrating on intricate areas (such as low-power zones with buildings) where radio propagation is difficult to predict, we focus on simpler areas (such as high-power zones without buildings) where radio propagation principles are more straightforward. By using radio propagation models to calculate the RSS values of points in these simpler areas, we can compare these calculated values with those in the radio map estimated by DDIM. We then consider the radio map whose RSS values closely align with those predicted by the propagation models as a high-quality radio map.
Specifically, during the election block, we consider areas pertaining to high-power zones situated in close proximity to transmitters as simpler areas. Given that the vicinity of the transmitter closely approximates the ideal conditions for electromagnetic wave propagation, the RSS distribution within this area is expected to adhere to the Friis Transmission Formula [26]:
@@
where $P_{t}$ is the power fed into the transmitting antenna at its input terminals, $P_{r}$ is power available at the output terminals of the receiving antenna, $A_{r}$ is the effective area of the receiving antenna, $A_{t}$ is the effective area of the transmitting antenna, $d$ is the distance between antennas, and $\lambda$ is the wavelength.
The Friis Transmission Formula pertains to the scenario involving two single-port antennas and is predicated on several key assumptions, including polarization matching, a sufficiently large distance $d$ (satisfying the far-field condition), and the reciprocity of the transmitting antenna. In the election block of our work, we employ a generalized version of the Friis Transmission Formula that is applicable to multiport antenna arrays [27]:
@@
where $\mathbf{y}_{\text{ii}}$ is the self-admittance of the input port, $\mathbf{y}_{\text{oi}}$ is the mutual admittance from the input port to the output port, $\mathbf{y}_{\text{oo}}$ is the self-admittance of the output port. The inner admittance marix $\mathbf{y}_{\text{si}}$ and the open circuit voltage $\mathbf{v}_{si}=\mathbf{v}_{i}+\mathbf{y}_{si}^{-1}\mathbf{i}_{i}$ is connected at the input port, and $\mathbf{x}=\left(\mathbf{y}_{ii}+\mathbf{y}_{si}\right)^{-1}\mathbf{y}_{si}\mathbf{v}_{si}$.
@@
Although information of transmitters are not inputs of our studied radio map estimation problem, at the beginning of the election block, we already get a candidate set of radio maps from DDIM and it is easy to learn the information of transmitters from each map. By leveraging the aforementioned principles of radio propagation in conjunction with the transmitter information we have learned, we can predict the RSS distribution, denoted as $\text{RSS}_{phy}(O,d,\theta)$, in the vicinity of the transmitters using polar coordinates. Here, $O$ represents the center of transmitters and the polar axis aligns with the x-axis of the original data.
To find the best map in the candidate set of maps generated by DDIM, we concentrate on the circular regions centered around the transmitters and compute the corresponding RSS values. We then compare these calculated RSS values with those at the same locations in each candidate map. The quality of a map candidate is determined by the extent of similarity, which we quantify using MSE. Furthermore, we emit probe rays from the transmitters and examine the variations in RSS as the rays propagate until they intersect with buildings. The credibility of our results is evaluated by assessing whether the observed RSS behavior conforms to the smooth attenuation pattern predicted by physical laws of radio propagation.
Fig. 9 gives an simulated example (the specific settings of simulations will be presented later in Section V) to illustrate how the election block uses laws of radio propagation to find best radio map from the set of map candidates. Fig. 9a is the true radio map with three transmitters. Both Fig. 9b and Fig. 9c are estimated radio maps, with Fig. 9b being better than Fig. 9c. In Fig. 9d, we compare the RSS values within a circular region centered around the first transmitter across different maps. It is evident that the RSS values obtained from the better map (blue results) closely match those of the true map (green results), while the RSS values from the worse map (red results) deviate significantly. Similar patterns are observed in Fig. 9e and Fig. 9f, which correspond to the circular regions centered around the second and third transmitters, respectively. Fig. 9g displays the changes in RSS values along rays emitted from the first transmitter. It is clear that the changes in RSS values in the better map (blue results) are much closer to zero compared to those in the worse map (red results). Similar observations can be found in Fig. 9h and Fig. 9i, respectively.
@@
where $n$ is the number of transmitters, $\alpha$ is a hyperparameter, and $\sigma$ is a threshold of volatility intensity. For the election block, we calculate MSE between $\mathbf{x}_{a}$ and $\mathbf{x}_{zi}$ with correction and $\text{rate}_{phy}$ to find the best map $\mathbf{x}_{z}$ when the sampling rate is ultra-low. Specifically, we define

### v3

@@ IV WiFi-Diffusion: Design Details
WiFi-Diffusion comprises three blocks: a boost block, a generation block, and an election block. In this section, we describe the design details of each WiFi-Diffusion block.
@@ IV-A Boost block
The objective of this block is to improve the quality of fine-grained radio maps produced by DDIM in the subsequent generation block. In this block, we design AttUnet, a UNet model with self-attention, to pinpoint “critical points” within the region $\mathcal{Y}$. AttUnet predicts RSS values at these critical points based on prior information, including the layout of buildings, the law of radio propagation, and RSS values of known radio samples. These predictions are then provided to DDIM. Although DDIM, as a diffusion model, is adept at focusing on important features, it lacks a deep understanding of the underlying physical laws. By incorporating AttUnet’s predictions that are aware of physical laws, DDIM’s performance can be improved. Furthermore, AttUnet enhances DDIM by providing it with RSS information of additional radio samples (i.e., with RSS values predicted for the identified critical points). Using both identified critical points and limited collected samples, DDIM can generate fine-grained radio maps of higher quality compared to those maps generated solely from the limited collected samples.
It should be clear that DDIM can more easily predict the propagation of high-power radios in an environment without buildings. In contrast, predicting the propagation of low-power radios in a complex environment with buildings poses a greater challenge to DDIM. We consider the complex environment with buildings as the area containing points that are critical to DDIM.
Figure 4: Architecture of AttUnet used in the boost block in WiFi-Diffusion.
@@
The architecture of AttUnet is illustrated in Fig. 4, and the parameters of AttUnet are detailed in Table II. Specifically, AttUnet first generates a rough estimate of the fine-grained radio map. Then, we consider critical points to be located in areas where there is a localized, abrupt change in the RSS or a sudden variation in the rate of change in the RSS on the map. These phenomena are typically attributed to intricate interactions between electromagnetic waves and buildings, as well as other waves. To identify these critical points, we shift a local window in various directions to analyze the area within the window. If this shifting results in substantial changes in the intensity of RSS, we consider the corresponding point to be critical.
@@
where $\mathbf{x}_{a}(x,y)$ is the RSS of the point located at $(x,y)$ in the map $\mathbf{x}_{a}$, and $w(x,y)$ is a Gaussian window function which could effectively remove the noise in $\mathbf{x}_{a}$ and is suitable for smoothing $\mathbf{x}_{a}$ with continuity due to its characteristics of high center weights and low edge weights. Local quadratic approximation of $E(u,v)$ in the neighborhood of $(0,0)$ is given by the second-order Taylor expansion:
@@
Given a point $(x,y)$, we first calculate $\mathbf{M}$ and then compute $\lambda_{1}$ and $\lambda_{2}$. If $\lambda_{1}$ and $\lambda_{2}$ are large, the RSS change of the map $\mathbf{x}_{a}$ at this point is abrupt, and we consider this point as a critical point.
Overall, in the boost block, we first use AttUnet to generate a preliminary fine-grained radio map from the collected known radio samples and buildings layout, and then identify critical points in the map from the perspective of laws of radio propagation. At the end of the boost block, we convey the predicted radio samples (identified critical points as well as corresponding predicted RSS values) and the known samples collected by the sensors to DDIM. To further enhance the performance of DDIM, we also randomly select a specific number of points in the map generated by AttUnet and transmit their predicted RSS values to DDIM. Similarly to $\mathbf{S}_{k}$, we define $\mathbf{ES}_{k}$ as a $H\times W$ matrix, where its element at the index $(h_{s},w_{s})$, i.e., $\mathbf{ES}_{k}(h_{s},w_{s})$, represents the RSS value at that grid point if it is a critical point, a randomly selected point, or a point where a radio sample has been collected; otherwise, it is 0. Fig. 5a gives an example of $\mathbf{S}_{k}$ and Fig. 5c gives an example of $\mathbf{ES}_{k}$. Fig. 5b gives an example of a radio map, where there are original points corresponding to known radio samples, critical points identified in the boost block, and randomly selected points in the map. $\mathbf{ES}_{k}$ is the output of the boost block and is the input of the subsequent generation block.
Figure 5: Different kinds of points that are the outputs of the boost block in WiFi-Diffusion.
@@ IV-B Generation block
The objective of this block is to generate a diverse set of fine-grained radio map candidates. We employ DDIM to achieve this goal.
We consider the task of fine-grained radio map estimation as a generative problem, in which a fine-grained radio map is regarded as a sample drawn from a $(H\times W)$-dimensional distribution. By learning this distribution from available radio map datasets, we can effectively tackle the challenge posed by ultra-low sampling rate and produce fine-grained, high-quality radio maps even when the number of known radio samples in the map to be estimated is extremely small.
We assume that $\mathbf{x}_{0}$, which is the fine-grained radio map to be estimated, follows a certain conditional probability distribution $\mathcal{X}_{c}$ given $\mathbf{B}$ and $\mathbf{ES}_{k}$. We use the Denoising Diffusion Probabilistic Model (DDPM) [20] to learn $\mathcal{X}_{c}$. The key idea of DDPM involves transforming $\mathcal{X}_{c}$ into a standard normal distribution $\mathcal{N}(\mathbf{0},\mathbf{I})$, and then training a neural network to model the reverse process, thereby establishing a mapping from $\mathcal{N}(\mathbf{0},\mathbf{I})$ back to $\mathcal{X}_{c}$. With DDPM, we can get sample values of $\mathcal{X}_{c}$ by sampling in $\mathcal{N}(\mathbf{0},\mathbf{I})$.
Specifically, DDPM consists of two processes: a forward process (diffusion process) and a reverse process (inverse diffusion process).
@@
where $\beta_{t}$ is the hyperparameter chosen before model training. Using the chain rule and the Markov property of probability, we denote the joint distribution of $\mathbf{x}_{1},\mathbf{x}_{2}\dots\mathbf{x}_{t}$ conditional on $\mathbf{x}_{0}$ as:
@@
where $\epsilon\sim\mathcal{N}(\mathbf{0},\mathbf{I})$. Let $\alpha_{t}:=1-\beta_{t},\bar{\alpha}_{t}:=\prod_{s=0}^{t}\alpha_{s}$. We can marginalize the above joint distribution to give:
@@
Define condition $\{\mathbf{B},\mathbf{ES}_{k}\}$ as $c$, we have:
@@
During inference, our aim is to generate a diverse set of results (fine-grained radio maps) from different noises sampled in $\mathcal{N}(\mathbf{0},\mathbf{I})$. A significant limitation of DDPM is the substantial number of iterations required to infer and generate high-quality results. To address this issue, we use DDIM [24] for inference. DDIM markedly reduces the inference time while preserving the high quality of the generated results. Furthermore, DDIM can be seamlessly applied after model training, allowing for a straightforward conversion of a DDPM model to a DDIM model without the need for retraining.
@@
where $\sigma_{t}^{2}=\eta\cdot\sqrt{(1-\alpha_{t-1})/(1-\alpha_{t})}\sqrt{(1-\alpha_{t}/\alpha_{t-1})}$. If $\eta=1$, the DDIM method is the same as DDPM. If $\eta=0$, the DDIM method solves the ordinary differential equations.
The training and inference of the generation block are illustrated in Fig. 6. At the end of the generation block, we get a diverse set of fine-grained radio map candidates generated by DDIM from $\mathbf{B}$ and $\mathbf{ES}_{k}$.
@@ IV-C Election block
The objective of this block is to identify the best map in the diverse set of fine-grained radio map candidates generated by the generation block. We leverage the radio propagation model as a guide to identify the map that adheres most closely to the propagation model in the set and consider it the best map.
@@
As discussed in the previous block and illustrated by Fig. 7, the diffusion model DDIM can build a mapping from the noisy distribution $\mathcal{N}(\mathbf{0},\mathbf{I})$ to the conditional distribution $\mathcal{X}_{c}$ under condition $c$:
@@
where $\mathbf{x}_{zi}\in\mathcal{X}_{c}$ and $\mathbf{z}_{i}\in\mathcal{N}(\mathbf{0},\mathbf{I})$ with $i\in\{1,2,\cdots,m\}$ and $m$ being the number of map candidates generated by DDIM.
The problem of radio map estimation requires the result $\mathbf{x}_{zi}$ to be close to the true radio map $\mathbf{x}_{0}$. Therefore, we generate $\{\mathbf{x}_{z1},\mathbf{x}_{z2},\cdots,\mathbf{x}_{zm}\}$ from $\{\mathbf{z}_{1},\mathbf{z}_{2},\cdots,\mathbf{z}_{m}\}$, and select $\mathbf{x}_{z}\in\{\mathbf{x}_{z1},\mathbf{x}_{z2},\cdots,\mathbf{x}_{zm}\}$ that is closest to $\mathbf{x}_{0}$ as output. Considering that the true map $\mathbf{x}_{0}$ is unknown, intuitively, we plan to use the Mean-Square Error (MSE) loss between the RSS values of known radio samples and the RSS values predicted by DDIM to approximate the distance between $\mathbf{x}_{zi}$ and $\mathbf{x}_{0}$, i.e.,
@@
where $\mathbf{mask}_{k}$ denotes a mask matrix consisting entirely of zeros, except for the element located at index $(h_{s},w_{s})$ which corresponds to a non-zero element in $\mathbf{S}_{k}$. From the above equation, we have
@@
However, we note that $\text{spMSE}(\mathbf{x}_{zi},\mathbf{S}_{k})$ does not approximate $\text{MSE}(\mathbf{x}_{zi},\mathbf{x}_{0})$ well in practice, mainly due to the ultra-low sampling rate (because almost all elements in $\mathbf{S}_{k}$ are zero). Recall that in the boost block, AttUnet generates a radio map $\mathbf{x}_{a}$. Here we use $\mathbf{x}_{a}$ to help select the best map. We introduce a correction factor $res$ as follows
@@
where the value of $res$ can be derived from different methods of analysis. ResNet [16] has been widely used in the past few years in feature extraction applications. In WiFi-Diffusion, we use it to predict $res$. We do this by providing $\mathbf{x}_{a}$, $\mathbf{x}_{zi}$, the Laplacian of Gaussain (LoG) operator processing information (edge detection), and high power maps of $x_{zi}$ (an illustrative example is shown in Fig. 8) to ResNet to learn the correction factor $res$.
@@
In addition to the MSE, we also incorporate a radio propagation model as a guide to select the best map during the election block. It is important to note that obtaining an exact radio propagation model for the radio map to be estimated is almost impossible, particularly in complex environments. Instead of focusing on intricate areas (such as low-power zones with buildings) where radio propagation is difficult to predict, we focus on simpler areas (such as high-power zones without buildings) where radio propagation is more straightforward. Using the radio propagation model to calculate the RSS values of the points in these simpler areas, we can compare these calculated values with those estimated by the radio map given by DDIM. We then consider the radio map whose estimated RSS values closely align with those calculated by the propagation model as a high-quality radio map.
Specifically, during the election block, we consider areas pertaining to high-power zones located in close proximity to transmitters as simpler areas. Given that the vicinity of the transmitter closely approximates the ideal conditions for radio propagation, the RSS distribution within this area is expected to adhere to the Friis Transmission Formula [25]:
@@
where $P_{t}$ is the power fed into the transmitting antenna at its input terminals, $P_{r}$ is the power available at the output terminals of the receiving antenna, $A_{r}$ is the effective area of the receiving antenna, $A_{t}$ is the effective area of the transmitting antenna, $d$ is the distance between antennas, and $\lambda$ is the wavelength.
The Friis Transmission Formula pertains to the scenario involving two single-port antennas and is predicated on several key assumptions, including polarization matching, a sufficiently large distance $d$ (satisfying the far-field condition), and the reciprocity of the transmitting antenna. In our election block, we employ a generalized version of the Friis Transmission Formula that is applicable to multiport antenna arrays [26]:
@@
where $\mathbf{y}_{\text{ii}}$ is the self-admittance of the input port, $\mathbf{y}_{\text{oi}}$ is the mutual admittance from the input port to the output port, and $\mathbf{y}_{\text{oo}}$ is the self-admittance of the output port. The inner admittance matrix $\mathbf{y}_{\text{si}}$ and the open circuit voltage $\mathbf{v}_{si}=\mathbf{v}_{i}+\mathbf{y}_{si}^{-1}\mathbf{i}_{i}$ are connected at the input port, and $\mathbf{x}=\left(\mathbf{y}_{ii}+\mathbf{y}_{si}\right)^{-1}\mathbf{y}_{si}\mathbf{v}_{si}$.
@@
Although the information of transmitters is not input of our studied radio map estimation problem, at the beginning of the election block, we already get a candidate set of fine-grained radio maps from DDIM. It is easy to learn the information about the transmitters from each map in the candidate set. Using the aforementioned principles of radio propagation in conjunction with the transmitter information we have learned, we can predict the RSS distribution, denoted as $\text{RSS}_{phy}(O,d,\theta)$, in the vicinity of the transmitters using polar coordinates. Here, $O$ represents the centers of transmitters, and the polar axis aligns with the x-axis of the original data.
To find the best map in the candidate set of maps generated by DDIM, we focus on the circular regions centered around the transmitters and compute the corresponding RSS values. We then compare these calculated RSS values with those at the same locations in each candidate map. The quality of a map candidate is determined by the degree of similarity, which we quantify using MSE. Furthermore, we emit probe rays from the transmitters and examine the variations in RSS as the rays propagate until they intersect with buildings. The credibility of our results is evaluated by assessing whether the observed RSS behavior conforms to the smooth attenuation pattern predicted by the physical laws of radio propagation.
Fig. 9 gives a simulated example (the specific settings of simulations will be presented later in Section V) to illustrate how the election block uses laws of radio propagation to find the best radio map from the set of map candidates. Fig. 9a is the true (ground truth) radio map with three transmitters. Both Fig. 9b and Fig. 9c are estimated radio maps, with Fig. 9b better than Fig. 9c. In Fig. 9d, we compare the RSS values within a circular region centered around the first transmitter on different maps. It is evident that the RSS values obtained from the better map (blue results) closely match those of the true map (green results), while the RSS values from the worse map (red results) deviate significantly. Similar patterns are observed in Fig. 9e and Fig. 9f, which correspond to the circular regions centered around the second and third transmitters, respectively. Fig. 9g displays the changes in RSS values along the rays emitted from the first transmitter. It is clear that the changes in RSS values in the better map (blue results) are much closer to zero compared to those in the worse map (red results). Similar observations can be found in Fig. 9h and Fig. 9i, respectively.
@@
where $n$ is the number of transmitters, $\alpha$ is a hyperparameter, and $\sigma$ is a threshold of volatility intensity. In the election block, we calculate the MSE between $\mathbf{x}_{a}$ and $\mathbf{x}_{zi}$, as well as $\text{rate}_{phy}$ to find the best map $\mathbf{x}_{z}$. Specifically, we define